OpenStudy (anonymous):

Write log((yx^2)/(sqr root z)) in terms of log x, log y, and log z

ganeshie8 (ganeshie8):

use : log(x/y) = log(x) - log(y) log(xy) = log(x) + log(y) log(x^n) = nlog(x)

ganeshie8 (ganeshie8):

what answer you got finally ?

OpenStudy (anonymous):

I got it to 2logx + logy - 1/2 log z and so now I'm just putting it into the terms of x y and z

ganeshie8 (ganeshie8):

great.. this is the answer right ?